Suppose that the production function is given as follows:

TPL = 10L + 5L2 + L3

Find the total product, Marginal product and average product when L = 5. 


1
Expert's answer
2021-05-06T14:04:10-0400

"TP_L = 10L + 5L^2 + L^3"


"MP_L=\frac {dTP_L} {dL} =10+10L+3L^2"


"AP_L=\frac {TP_L} {L} =10+5L+L^2"


When L=5;

"TP_L=" "10(5) + 5(5)^2 + (5)^3"


"TP_L=50+125+125"


"TP_L=300"



"MP_L=10+10(5)+3(5)^2"


"MP_L=10+50+75"


"MP_L=135"



"AP_L=10+5(5)+(5)^2"


"AP_L=60"